OpenStudy (alisonxfam):

What affect does h = 2 have on the function f(x) = log(x)? The graph moves 2 units to the right. The graph moves 2 units to up. The graph moves 2 units to the left. The graph is stretched by a factor of |2|.

OpenStudy (alisonxfam):

would my answer be The graph moves 2 units to up.

OpenStudy (alisonxfam):

@DanJS

OpenStudy (danjs):

no, remember the h value is the horizontal shift amount. here the graph shifts horizontal by an amount 'h' --- f(x) = log( x-h ) = log(x - 2)